OpenStudy (anonymous):

Solve each system of equations by using elimination. 3p+9q=6 5p-5q=30

OpenStudy (anonymous):

Think about eliminating the -5q with the 9q How are you going to do it ?

OpenStudy (anonymous):

Im not sure how to solve that, im used to it with opposites there already. for example if it was a 9q and a -9q i could solve it

OpenStudy (anonymous):

Yes So you have to Multiply one of the two equations by a number which could make the -9q and of course you will muliply the one which contains the -5q

OpenStudy (anonymous):

but, how are we going to get this number

OpenStudy (anonymous):

AM i multiplying it by -1

OpenStudy (anonymous):

No you have to look for a factor which relates the -5 and the -9 in other words which number will convert the -5 to -9

OpenStudy (anonymous):

1.8

OpenStudy (anonymous):

yes So do it

OpenStudy (anonymous):

I got 3p + 9q=6 and 9p-9q=54. Is this right?

OpenStudy (anonymous):

After doing it you should have the two equations: \[3p+9q=6\] 9p-9q =54

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

Now Add the two equations

OpenStudy (anonymous):

i got p = 5

OpenStudy (anonymous):

You notice that -9q and 9q is the same number but different signs,so upon adding they will be eliminated together

OpenStudy (anonymous):

Exactly, Great Job.well Done this is the adding by eliminating

OpenStudy (anonymous):

ANd i plug the 5 in right?

jhonyy9 (jhonyy9):

so you know that 5p-5q = 30 and again that p=5 than substitute the 5 in place of p and calculi the value of q ok. ?